OpenStudy (anonymous):

Find the derivative: sin(xcosx) so the answer is after being simplified (cosx-xsinx)cos(xcosx) I have no idea. how.

OpenStudy (anonymous):

this is chain rule

OpenStudy (anonymous):

\[\frac{dy}{dx}=\frac{dy}{du}\frac{du}{dx}\] where your y=sin(u) and your u=xcos(x)

OpenStudy (anonymous):

ok so: sin(xcosx)(-sinx)(1)+cos(xcosx)

OpenStudy (anonymous):

why is that not the answer I am getting :(

OpenStudy (anonymous):

alright so derivative of sin(u) = cos(u) so \[\frac{dy}{du}=cos(u)\] next do the product rule with your u=xcos(x) u'=vw'+wv'=-xsin(x)+cos(x)

OpenStudy (anonymous):

so you end up with cos(xcos(x)(-xsin(x)+cos(x))

OpenStudy (anonymous):

which is the same as (cos(x)-xsin(x))cos(xcos(x))

OpenStudy (anonymous):

does this make sense?

OpenStudy (anonymous):

hi